### v1

@@ Appendix B CGP-CDE Details
For our Bayesian causal model, we use the Gaussian process conditional density estimator [Dutordoir et al., 2018]. We use a sum of commonly used kernels [Rasmussen, 2003, Ch. 5], as outlined in section B.1. We also apply variational inference to optimise the lower bound, as is standard for latent variable Gaussian process models [Dutordoir et al., 2018, Titsias and Lawrence, 2010] and use inducing points with stochastic variational inference to improve data scalability [Hensman et al., 2013]. We outline the variational inference in section B.2 and detail our optimisation schedule in section B.3.
B.1 CGP-CDE Kernel
@@
The ultimate aim of the approach is to find the causal graph that maximises the posterior probability. The choice of kernels will be problem dependent and it is possible to carry out Bayesian model selection over the kernels themselves [Rasmussen, 2003, Ch. 5]. To allow for good fits of a range of datasets, we use a sum of kernels that can express functions with a range of lengthscales and roughness well.
@@
where each of the kernels is defined below [Williams and Rasmussen, 2006, Ch. 4]. For all the kernels, $\theta_{ii}$ is the hyperparameter for the latent dimension.
@@
$k_{\text{rq}}$ is the rational quadratic kernel, which is equivalent to the sum of many squared exponential kernels with different precision hyperparameters [Williams and Rasmussen, 2006, Ch. 4],
@@
B.2 CGP-CDE Lower Bound
We use a variant on the GP-CDE [Dutordoir et al., 2018] to define our causal model and define a prior over distributions. Intractability in these models is commonly solved using variational inference [Lalchand et al., 2022, Dutordoir et al., 2018].
@@ B.2 CGP-CDE Lower Bound
The presence of the latent variable allows learning of non-Gaussian and heteroscedastic noise [Dutordoir et al., 2018]. A Gaussian process prior is placed on $\mathbf{f}_{i}$,
@@
The addition of inducing points $\mathbf{u}_{i}$, with corresponding inducing inputs $\boldsymbol{Z}_{i}$, helps both with the intractability of $\log p(\mathbf{x}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)},\boldsymbol{\Lambda}_{i},\boldsymbol{\phi}_{i})$ and scaling to large datasets [Titsias, 2009, Titsias and Lawrence, 2010]. We can write the joint between $\mathbf{f}$ and $\mathbf{u}_{i}$ as:
@@
where $p(\mathbf{u}_{i})=\mathcal{N}(\boldsymbol{0},K_{i,\mathbf{u}\mathbf{u}})$. This is still not tractable, so we use variational inference to define an evidence lower bound (ELBO) to the marginal likelihood [Titsias, 2009]. This is done by introducing a variational distribution $q(\mathbf{f}_{i},\mathbf{w}_{i},\mathbf{u}_{i})$ and rewriting the marginal likelihood as (suppressing conditioning terms for neatness):
@@
The variational distribution takes the form [Titsias and Lawrence, 2010, Dutordoir et al., 2018]:
@@
where $q(\mathbf{u}_{i})=\mathbf{N}(\mathbf{u}_{i}|\boldsymbol{m}_{u,i},\boldsymbol{S}_{u,i})$ is the Gaussian variational distribution of inducing points $\mathbf{u}_{i}$ and $q(\mathbf{w}_{i})=\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ is the variational distribution of $\mathbf{w}_{i}$.
Rearranging and using Jensen’s inequality, we can then get the lower bound for variable $d$ as [Dutordoir et al., 2018]
@@
By incorporating the optimal Gaussian variational distribution $q(\mathbf{u})$, which can be calculated in closed form, it is possible to collapse the bound by integrating out the inducing variables [Titsias, 2009, Titsias and Lawrence, 2010]. However, the computational complexity of this approach is $\mathcal{O}(NM^{2}D)$, meaning for large values of $N$ training becomes infeasible [Lalchand et al., 2022]. To allow our method to scale to large datasets, we follow the approach of Hensman et al. [2013] and keep $q(\mathbf{u})$ uncollapsed so that stochastic variational inference can be used. This means the data can be minibatched in training, reducing the computational complexity to $\mathcal{O}(M^{3}D)$. To further improve computational efficiency, we use an encoder to learn a function $h_{q,i}:(\mathbf{x}_{\text{batch}})\mapsto(\mu_{i},\Sigma_{i})$ for the variational distribution $q(\mathbf{w}_{i})=\mathcal{N}(\mu_{i},\Sigma_{i})$, where $\mathbf{x}_{\text{batch}}\in\mathbb{R}^{b\times D}$ is a batch of data points, $\mu_{i}\in\mathbb{R}^{b}$ and $\Sigma_{i}\in\mathbb{R}^{b}$ for batch size $b$ [Dutordoir et al., 2018]. This is more efficient than the alternative of learning $N$ parameters. We use a multi-layer perceptron for the encoder and list the hyperparameters used in table 2.
Table 2: Hyperparameters for the variational encoder $h_{q,i}:(\mathbf{x}_{\text{batch}})\mapsto(\mu_{i},\Sigma_{i})$.
@@
We follow common practice and maximise the lower bound with respect to the hyperparameters $\boldsymbol{\sigma}$ and $\boldsymbol{\phi}$ as well [Rasmussen, 2003] (along with $\boldsymbol{\theta}$ and $q$). This is justified based on the fact that the posterior for these hyperparameters tends to be highly peaked in practice [MacKay, 1999]. Thus, a Laplace approximation can be approximated by the maximum value.
B.3 Optimisation schedule
The final score, optimised using the augmented Lagrange method [Zheng et al., 2018, Lachapelle et al., 2019], is
$\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi}):=\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})+\log p(\boldsymbol{\theta})-\alpha_{t}|h(\mathbf{A}_{\boldsymbol{\theta}})|^{2}-\frac{\rho_{t}}{2}h(\mathbf{A}_{\boldsymbol{\theta}}),$ (24)
where $t$ is the index of the subproblem and $\alpha_{t},\rho_{t}$ are hyperparameters that are increased after each subproblem. We consider a subproblem "solved" when $\mathcal{L}_{\text{ELBO}}$ stops increasing. We follow Nemirovsky [1999] and previous work [Lachapelle et al., 2019, Yu et al., 2019] in the update schedule for $\alpha_{t}$ and $\rho_{t}$. We initialise $\rho_{0}=0$ and $\alpha_{0}$ heuristically such that the value of $\alpha_{0}|h(\mathbf{A}_{\boldsymbol{\theta}})|^{2}$ is a fraction of the scale of $\mathcal{L}_{\text{ELBO}}$. The scale of $\mathcal{L}_{\text{ELBO}}$ can be found by simply running the GP-CDEs without the causal component. After each subproblem we update
$\rho_{t+1}$ $\leftarrow\rho_{t}+\alpha_{t}h(\mathbf{A}_{\boldsymbol{\theta}_{t}}),$ (25)
$\alpha_{t+1}$$\leftarrow\begin{cases}\nu\alpha_{t}\text{ if }h(\mathbf{A}_{\boldsymbol{\theta}_{t}})>\gamma h(\mathbf{A}_{\boldsymbol{\theta}_{t-1}})\\ \alpha_{t}\text{ Otherwise}\end{cases}$ (26)
where we set $\nu=10.0$ and $\gamma=0.9$. Optimisation is stopped when $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})<\epsilon$.
As Gaussian process models with latent variables suffer from local minima, we found that breaking up training into phases aided the optimisation. We also outline other choices that helped.
Warm up phase:
This phase of training was the first subproblem ($t=0$) of eq. 24. As the hyperparameters of kernels of Gaussian processes can take a while to converge to a solution, we found making this phase of training longer than necessary helpful. We achieved this by ensuring training took at least $T_{0}$ iterations before checking if the first subproblem was solved. The value of $T_{0}$ will depend on the problem at hand, but for our experiments, $T_{0}=30,000$ sufficed.
@@ B.3 Optimisation schedule
This is the main part of the training, where the loss described in eq. 24 is optimised. This phase is terminated when $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})<\epsilon$.
Cool down phase:
The value of the ELBO found at the end of training (when $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})<\epsilon$) is highly dependent on how many subproblems the method had to solve. To allow for the comparison of ELBO values across runs (see section B.4) we continue training the found acyclic DAG for $T_{f}$ iterations. This ensured that the found ELBO approximates the ELBO of the final resultant DAG. As some hyperparameters may have been pushed to extreme values by the acyclic threshold if competing with an anticausal edge, we reinitialise the hyperparameters of the edges that are still active to their initial value. This reduces the number of iterations needed for the cool down phase. We found that $T_{f}=30,000$ cool down iterations sufficed here.
Thresholding:
The values of the hyperparameters can remain very low but the function learnt will still be constant. Following previous work [Lachapelle et al., 2019], to aid with learning the adjacency matrix, we threshold values of hyperparameters to $10^{-15}$ if they become lower than $10^{-4}$ after the warm up phase.
Cool down thresholding:
Before the start of the cool-down phase, as the values of the hyperparameters sometimes are not exactly 00 due to numerical precision, we find the adjacency matrix with an extra thresholding step. Following Lachapelle et al. [2019], we threshold weights to 00, starting with the lowest weight, until the adjacency is exactly acyclic ($h(\mathbf{A}_{\boldsymbol{\theta}_{t}})=0$).
@@
To extract the causal graph at the end of training, we remove any edges where the hyperparameters do not encode correlation. To do this, we threshold the linear variances to $10^{-4}$ and the graph parameters to $0.05$, which corresponds to a lengthscale of $20$.
B.4 Random Restarts
The main principle behind our method is to find the graph that maximises the marginal likelihood (or a lower bound to it). Maximising the ELBO with respect to the hyperparameters $\boldsymbol{\Lambda}$ in Gaussian process models is known to suffer from local optima issues [Rasmussen, 2003, Ch. 5]. The final hyperparameters found, and hence the adjacency matrix, can be dependent on the initialisation of the hyperparameters. To mitigate this, we use a widely adopted technique of performing random restarts [Dhir et al., 2024]. We optimise the loss $\mathcal{L}$ starting from multiple initialisations (this can be done in parallel) $N_{r}$ times. Then we pick the graph that achieves the highest $\mathcal{L}_{\text{ELBO}}$ as the candidate for the most likely graph. Adding more random restarts for the CGP-CDE should improve performance, as shown for the DGP-CDE section C.1.
Input: Data $\mathbf{X}$, number of random restarts $N_{r}$, $\alpha_{0}$, $\rho_{0}$, initial $\boldsymbol{\Lambda}=\{\boldsymbol{\theta},\boldsymbol{\sigma}\}$
@@ B.4 Random Restarts
Initialise empty list $graphscore$for $i=1$ to $N_{r}$ do
@@
             Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (eq. 24)
@@
      $t\leftarrow 0$while $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})>\epsilon$ do
             if  $\mathcal{L}$ is increasing  then
                   Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (eq. 24)
            else
                   $t\leftarrow t+1$$\alpha_{t}\leftarrow\alpha_{t+1}$$\rho_{t}\leftarrow\rho_{t+1}$
@@
      Set $\alpha_{t},\rho_{t}$ to 00  Threshold $\mathbf{A}_{\boldsymbol{\theta}_{t}}$ starting from lowest weight until $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})=0$for $k=1$ to $T_{f}$ do
            Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (eq. 24)
@@
      Append $(\mathbf{A}_{\boldsymbol{\theta}},\mathcal{L})$ to $graphscore$.
@@
$\mathbf{A}^{*}\leftarrow\mathbf{A}_{\boldsymbol{\theta}}$ from $graphscore$ with the maximum $\mathcal{L}$.
@@
B.5 Hyperparameter Priors
As discussed in subsection 3.3, we can place priors on the graphs by placing priors on the hyperparameters $\boldsymbol{\theta}$ of the CGP-CDE. For our implementation, we place a prior on $\boldsymbol{\theta}$ that favours sparser graphs. Specifically, we use the Gamma prior $P(\boldsymbol{\theta})=\text{Gamma}(\eta,\beta)$, where $\eta$ is the shape parameter and $\beta$ is the rate parameter. For all experiments, we set $\eta=1$ and $\beta=10$. Our hyperparameters for the priors were heuristically chosen and give log probabilities that are a fraction of the rest of the loss.
B.6 Implementation Details

### v2

@@ Appendix C CGP-CDE Details
For our Bayesian causal model, we use the Gaussian process conditional density estimator (Dutordoir et al., 2018). We use a sum of commonly used kernels Rasmussen (2003, Ch. 5), as outlined in Section C.1. We also apply variational inference to optimise the lower bound, as is standard for latent variable Gaussian process models (Dutordoir et al., 2018; Titsias & Lawrence, 2010) and use inducing points with stochastic variational inference to improve data scalability (Hensman et al., 2013). We outline the variational inference in Section C.2 and detail our optimisation schedule in Section C.3.
C.1 CGP-CDE Kernel
@@
The ultimate aim of the approach is to find the causal graph that maximises the posterior probability. The choice of kernels will be problem dependent and it is possible to carry out Bayesian model selection over the kernels themselves (Rasmussen, 2003, Ch. 5). To allow for good fits of a range of datasets, we use a sum of kernels that can express functions with a range of lengthscales and roughness well.
@@
where each of the kernels is defined below (Williams & Rasmussen, 2006, Ch. 4). For all the kernels, $\theta_{ii}$ is the hyperparameter for the latent dimension.
@@
$k_{\text{rq}}$ is the rational quadratic kernel, which is equivalent to the sum of many squared exponential kernels with different precision hyperparameters (Williams & Rasmussen, 2006, Ch. 4),
@@
C.2 CGP-CDE Lower Bound
We use a variant on the GP-CDE (Dutordoir et al., 2018) to define our causal model and define a prior over distributions. Intractability in these models is commonly solved using variational inference (Lalchand et al., 2022; Dutordoir et al., 2018).
@@ C.2 CGP-CDE Lower Bound
The presence of the latent variable allows learning of non-Gaussian and heteroscedastic noise (Dutordoir et al., 2018). A Gaussian process prior is placed on $\mathbf{f}_{i}$,
@@
The addition of inducing points $\mathbf{u}_{i}$, with corresponding inducing inputs $\boldsymbol{Z}_{i}$, helps both with the intractability of $\log p(\mathbf{x}_{i}|\mathbf{X}_{\mathrm{PA}_{\mathcal{G}^{\mathbf{A}}}(i)},\boldsymbol{\Lambda}_{i},\boldsymbol{\phi}_{i})$ and scaling to large datasets (Titsias, 2009; Titsias & Lawrence, 2010). We can write the joint between $\mathbf{f}$ and $\mathbf{u}_{i}$ as:
@@
where $p(\mathbf{u}_{i})=\mathcal{N}(\boldsymbol{0},K_{i,\mathbf{u}\mathbf{u}})$. This is still not tractable, so we use variational inference to define an evidence lower bound (ELBO) to the marginal likelihood (Titsias, 2009). This is done by introducing a variational distribution $q(\mathbf{f}_{i},\mathbf{w}_{i},\mathbf{u}_{i})$ and rewriting the marginal likelihood as (suppressing conditioning terms for neatness):
@@
The variational distribution takes the form (Titsias & Lawrence, 2010; Dutordoir et al., 2018):
@@
where $q(\mathbf{u}_{i})=\mathcal{N}(\mathbf{u}_{i}|\boldsymbol{m}_{u,i},\boldsymbol{S}_{u,i})$ is the Gaussian variational distribution of inducing points $\mathbf{u}_{i}$ and $q(\mathbf{w}_{i})=\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ is the variational distribution of $\mathbf{w}_{i}$.
Rearranging and using Jensen’s inequality, we can then get the lower bound for variable $d$ as (Dutordoir et al., 2018)
@@
By incorporating the optimal Gaussian variational distribution $q(\mathbf{u})$, which can be calculated in closed form, it is possible to collapse the bound by integrating out the inducing variables (Titsias, 2009; Titsias & Lawrence, 2010). However, the computational complexity of this approach is $\mathcal{O}(NM^{2}D)$, meaning for large values of $N$ training becomes infeasible (Lalchand et al., 2022). To allow our method to scale to large datasets, we follow the approach of Hensman et al. (2013) and keep $q(\mathbf{u})$ uncollapsed so that stochastic variational inference can be used. This means the data can be mini-batched in training, reducing the computational complexity to $\mathcal{O}(M^{3}D)$. To further improve computational efficiency, we use an encoder to learn a function $g_{q,i}:(\mathbf{x}_{\text{batch}})\mapsto(\mu_{i},\Sigma_{i})$ for the variational distribution $q(\mathbf{w}_{i})=\mathcal{N}(\mu_{i},\Sigma_{i})$, where $\mathbf{x}_{\text{batch}}\in\mathbb{R}^{b\times D}$ is a batch of data points, $\mu_{i}\in\mathbb{R}^{b}$ and $\Sigma_{i}\in\mathbb{R}^{b}$ for batch size $b$ (Dutordoir et al., 2018). This is more efficient than the alternative of learning $N$ parameters. We use a multi-layer perceptron for the encoder and list the hyperparameters used in Table 1.
Table 1: Hyperparameters for the variational encoder $g_{q,i}:(\mathbf{x}_{\text{batch}})\mapsto(\mu_{i},\Sigma_{i})$.
@@
We follow common practice and maximise the lower bound with respect to the hyperparameters $\boldsymbol{\sigma}$ and $\boldsymbol{\phi}$ as well (Rasmussen, 2003) (along with $\boldsymbol{\theta}$ and $q$). This is justified based on the fact that the posterior for these hyperparameters tends to be highly peaked in practice (MacKay, 1999). Thus, a Laplace approximation can be approximated by the maximum value.
C.3 Optimisation schedule
We split the training schedule up into three steps. The warm-up phase allows the model to learn conditional independences present in the data. In the acyclic constraint phase, where a penalty method is used to enforce the DAG constraint, the model then learns the direction of any edges. The learnt DAG is then enforced, and the ELBO is optimised in the cool-down phase to find the ELBO of the final model, which can then be used for comparing random restarts.
Warm-up phase:
The warm-up phase allows the model to learn any conditional independences from the data. It also ensures the hyperparameters and variational parameters of the Gaussian process are at reasonable values at the beginning of the acyclic phase. Although the number of iterations needed will depend on the problem at hand, we found $T_{0}=25,000$ iterations for this phase sufficient to ensure convergence.
@@ C.3 Optimisation schedule
After the warm-up phase has concluded, a lot of the edges corresponding to conditional independences are switched off. However, for variables that are correlated, the causal direction has not been decided. This part of the training forces the constraint that the final graph has to be acyclic. The loss is
$\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi}):=\mathcal{L}_{\text{ELBO}}(q,\boldsymbol{\theta},\boldsymbol{\sigma},\boldsymbol{\phi})+\log p(\boldsymbol{\theta})-\gamma_{t}h(\mathbf{A}),$ (30)
is optimised, using the acyclic regulariser proposed by Nazaret et al. (2024). The weighting $\gamma_{t}$ of the acyclic penalty term $h(\mathbf{A})$ is increased linearly each epoch by $\rho=50$, starting from 00. As we use minibatching, one epoch is a full pass through the dataset. For some of the $50$ variable datasets, we increased this to $\rho=250$ to reduce the time required to train. This is because the unconstrained loss scales with the number of variables, so for the $50$ variable dataset, a higher $\rho$ ensures a higher ratio between the acyclic constraint and the rest of the loss, ensuring faster convergence.
We use $50$ power iterations to approximate the eigenvectors required to calculate the gradients of the acyclic constraint, as described in Section A.2. The acyclic constraint phase is terminated when $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})<\tau$ where $\tau=0.005$, or the number of iterations exceeds a maximum value of $T_{a}=50,000$.
Acyclic thresholding:
After the end of the acyclic phase, the values of the hyperparameters are sometimes not exactly 00 due to numerical precision, we find the adjacency matrix with an extra thresholding step. Following Lachapelle et al. (2019), we threshold weights to 00, starting with the lowest weight, until the adjacency is exactly acyclic ($h(\mathbf{A}_{\boldsymbol{\theta}_{t}})=0$).
Cool-down phase:
We found that random restarts helped tackle local optima in optimisation Section C.4. To allow for the comparison of ELBO values across restarts (see Section C.4) we continue training the found acyclic DAG for $T_{f}$ iterations. This ensured that the found ELBO approximates the ELBO of the final resultant DAG. As some hyperparameters may have been pushed to extreme values by the acyclic threshold if competing with an anti-causal edge, we reinitialise the hyperparameters of the edges that are still active to their initial value. This reduces the number of iterations needed for the cool-down phase. We found that $T_{f}=25,000$ cool-down iterations sufficed here.
@@
Finally, we perform an extra thresholding step to remove edges that do not encode correlations. As our data is normalised, we check the kernel evaluation of points $-3$ and $3$, which should bound $95\%$ of the data. If the function is highly correlated between these points, it is effectively constant for $95\%$ of the data. We thus threshold the linear variances to $10^{-4}$ and the graph parameters (the rest of the kernels) to $0.05$, which corresponds to a lengthscale of $20$. These values ensure that points between $-3,3$ are highly correlated.
C.4 Random Restarts
The main principle behind our method is to find the graph that maximises the marginal likelihood (or a lower bound to it). Maximising the ELBO with respect to the hyperparameters $\boldsymbol{\Lambda}$ in Gaussian process models is known to suffer from local optima issues (Rasmussen, 2003, Ch. 5). The final hyperparameters found, and hence the adjacency matrix, can be dependent on the initialisation of the hyperparameters. To mitigate this, we use a widely adopted technique of performing random restarts (Dhir et al., 2024). We optimise the loss $\mathcal{L}$ starting from multiple initialisations (this can be done in parallel) $N_{r}$ times. Then we pick the graph that achieves the highest $\mathcal{L}_{\text{ELBO}}$ as the candidate for the most likely graph. Adding more random restarts for the CGP-CDE should improve performance, as shown for the DGP-CDE Section D.1.
Input: Data $\mathbf{X}$, number of random restarts $N_{r}$, acyclic penalty weighting update $\rho$, threshold $\tau$, acyclic penalty weighting $\gamma_{t}=0$ initial $\boldsymbol{\Lambda}=\{\boldsymbol{\theta},\boldsymbol{\sigma}\}$
@@ C.4 Random Restarts
Initialise empty list $graphscore$
for $i=1$ to $N_{r}$ do
@@
             Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (Equation 30)
@@
      $t\leftarrow 0$while $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})>\tau$ and $t<T_{a}$ do
             Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (Equation 30)  $t\leftarrow t+1$if  epoch ended then
                  $\gamma_{t}\leftarrow\gamma_{t}+\rho$
@@
Set $\gamma_{t}$ to 00  Threshold $\mathbf{A}_{\boldsymbol{\theta}_{t}}$ starting from lowest weight until $h(\mathbf{A}_{\boldsymbol{\theta}_{t}})=0$for $k=1$ to $T_{f}$ do
      Update $q,\boldsymbol{\Lambda}$ by maximising $\mathcal{L}$ (Equation 30)
@@
Append $(\mathbf{A}_{\boldsymbol{\theta}},\mathcal{L})$ to $graphscore$. $\mathbf{A}^{*}\leftarrow\mathbf{A}_{\boldsymbol{\theta}}$ from $graphscore$ with the maximum $\mathcal{L}$.
@@
C.5 Hyperparameter Priors
As discussed in Section 4.3, we can place priors on the graphs by placing priors on the hyperparameters $\boldsymbol{\theta}$ of the CGP-CDE. For our implementation, we place a prior on $\boldsymbol{\theta}$ that favours sparser graphs. Specifically, we use the Gamma prior $P(\boldsymbol{\theta})=\text{Gamma}(\eta,\beta)$, where $\eta$ is the shape parameter and $\beta$ is the rate parameter. For all experiments, we set $\eta=1$ and $\beta=10$. Our hyperparameters for the priors were heuristically chosen and give log probabilities that are a fraction of the rest of the loss.
C.6 Implementation Details